OpenStudy (anonymous):

Not sure of how to do this? Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3,-4)

OpenStudy (anonymous):

\[(y-y_{1})=m(x-x_{1})\] put the point and slope in this equation

OpenStudy (anonymous):

(3-y1)=-3(-4-x1) Huh?

OpenStudy (anonymous):

nope (x1,y1) is the point. x and y are variables

OpenStudy (anonymous):

Okay so like this (y-3)=-3(x--4)

OpenStudy (anonymous):

Thank you Greatly!! :)

hero (hero):

Also the method I've shown uses less steps and you only need to memorize one formula. All you need is y = mx + b 1). Substitute the given points and the slope to find b: -4 = -3(3) + b -4 = -9 + b 9 - 4 = b 5 = b 2). Re-write y = mx + b but only include the values of m and b: y = -3x + 5

hero (hero):

I think you get it now